California Fishing Company (CFC) is expected to pay a dividend next year of $50 per share. Future Dividends for CFC are expected to grow at a rate of 5% per year indefinitely. If an investor is currently willing to pay $500 each CFC share, what is the investor’s required return for this investment?
Answer #1

here mentioned that, it is expected dividend for next year.

Stock price = Dividend to be paid for the one year from now / (Required rate of return - Dividend growth)

⇒ $500 = $50 / (Required rate of return - 5%)

⇒ 500 × Required rate of return - (500 × 0.05) = 50

⇒ Required rate of return = 75/500 = 0.15

In this manner, the speculator's required rate of return is 15%.